OpenStudy (kittiwitti1):

A 3.00 kg block starts from rest at the top of a 30.0 degree incline and accelerates uniformly down the incline, moving 2.00 m in 1.50 s --- a. Find the magnitude of the acceleration of the block. b. Find the coefficient of kinetic friction between the block and the incline. c. Find the magnitude of the frictional force acting on the block. d. Find the speed of the block after it has slid a distance of 2.00 m.

OpenStudy (kittiwitti1):

I get b-c but I can't find a -_-

OpenStudy (anonymous):

For a) just use the first kinematic equation with initial velocity set to zero \[ d=\frac{1}{2}at^2\] d=2m t=1.5s b)  sum the forces along the ramp and perp to the ramp, with +x to the right \[ \sum F_x = F_f - mg \sin \theta = ma_{block}\] \[ F_f = ma_{block} + mg \sin \theta\] for c) find normal force \[ \sum F_y = N - mg \cos \theta = 0 \\ \ \ \hspace{120px} N=mg \cos \theta\] d) is another kinematic, use the fourth \[v_f^2=\cancel{v_i^2}^0 + 2ad\]

OpenStudy (anonymous):

I don't know why I was thinking Chemistry...

OpenStudy (anonymous):

Sorry, but I'm terrible at Physics, I;m good at Chemistry though, so if you need Chemistry help :P

OpenStudy (anonymous):

bueno o no?

OpenStudy (kittiwitti1):

Sorry, I had to take a break from the computer screen. Eyebrows hurt for some reason = n =;;

OpenStudy (kittiwitti1):

@bookworm00981 I'm pretty good with Chemistry too, but sure I'll consider it (;

OpenStudy (anonymous):

K :P

OpenStudy (kittiwitti1):

@AllTehMaffs Hold on, I've gotta look it over! D;

OpenStudy (kittiwitti1):

OMG you're amazing. Medal to you for sure! :3